OpenStudy (anonymous):

Can someone explain how ∫cos2xdx = (1/2)sin2x

OpenStudy (accessdenied):

try using u-substitution, with u = 2x. the only other thing to know is that the integral of cosine is sine. (because the derivative of sine is cosine)

OpenStudy (accessdenied):

and you can also confirm that it is true because the derivative of sin(2x)/2 = cos(2x)/2 * 2 = cos(2x)

OpenStudy (anonymous):

ohhhh I see it now thanks

OpenStudy (accessdenied):

No problem. :)